OpenStudy (anonymous):

find the solution to the system of equations by using either graphing or substitution. y = 6 - x and y = x -2 a. (2,4) b. (-4,2) c. (4,2) d. no solutions

OpenStudy (anonymous):

Okay so

OpenStudy (anonymous):

y=6−x y=x−2 Eliminate the equal sides of each equation and combine. 6−x=x−2 Since x contains the variable to solve for, move it to the left-hand side of the equation by subtracting x from both sides. 6−x−x=−2 y=x−2 Since −x and −x are like terms, subtract x from −x to get −2x. 6−2x=−2 y=x−2 Since 6 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 6 from both sides. −2x=−6−2 y=x−2 Subtract 2 from −6 to get −8. −2x=−8 y=x−2 Divide each term in the equation by −2. x=4 y=x−2 Replace all occurrences of x with the solution found by solving the last equation for x. In this case, the value substituted is 4. x=4 y=(4)−2

OpenStudy (anonymous):

x=4 y=4−2 Subtract 2 from 4 to get 2. x=4 y=2